OpenStudy (johnb):

Suppose I am a grandmother. My granddaughter is about as old in days as my daughter is in weeks. She is also as many months old as I am in years. Together, the three of us are 140 years old. How old am I?

OpenStudy (francesca_marie):

Suppose I am a grandmother. let a = grandma's age let d = daughter's age let g = granddaughter's age My granddaughter is about as old in days as my daughter is in weeks. 365g = 52d divide both sides by 52 Approx 7g = d She is also as many months old as I am in years. 12g = a Together, the three of us are 140 years old. a + d + g = 140 replace d with 7g, replace a with 12g 12g + 7g + g = 140 20g = 140 g = 140/20 g = 7 yrs old is the granddaughter : How old am I? a = 12(7) a = 84 yrs is grandma

OpenStudy (francesca_marie):

so shes 84

OpenStudy (francesca_marie):

It's kinda hard to explain